From mar@roe.ac.uk Tue Nov 5 14:11:53 2013 Date: Mon, 4 Nov 2013 16:54:21 +0000 From: Mike Read To: VDFS Science Archive Team -- Clive Davenhall , Dave Morris , Eckhard Sutorius , Mark Holliman , Nicholas Cross , Nigel Hambly , Rob Blake , Ross Collins , Stelios Voutsinas Cc: CCs for VDFS -- Alastair Edge , Andrew Lawrence , Bob Mann , Dante Minniti , James Dunlop , Jim Emerson , Keith Noddle , Lorenzo Rimoldini , Maria-Rosa Cioni , matt.jarvis@astro.ox.ac.uk, Mike Irwin , Nigel Metcalfe , Norman Gray , Phil Lucas , Richard McMahon at IOA , Simon Dye , Stephen Warren , Tom Kerr , Tom Shanks Subject: WFAU VDFS Science Archives progress meeting minutes, 1 November 2013 WFAU VDFS Science Archives progress meeting minutes, 1 November 2013 Present: RGM, MAR, RSC, RPB, NJC, ACD, ETWS, MSH Apologies: NCH, KTN, AL DoNM: 10am Friday 15 November 2013 in the Villas meeting room ACD in the chair (room is booked). Actions discharged this time: ---------------------------- ACTION: MAR/ETWS to investigate/fix the orientation of the N-E markings on the jpegs. ETWS confirms OSA and VSA jpegs have been fixed. ACTION: ACD, RGM to brief Stelios Voutsinas to arrange a meeting with Ken Rice. Done, the current access to HARPS N-data is adequate. ACTION: ACD to bring this item to ETWS' attention on his return. Done, in real time! This relates to VISTA v1.3.1 update. ACTION: MAR to produce an estimate of when the next OSA release is likely to be available. Done, the P90 ATLAS release would be made later that day! Actions partly discharged but continuing: ---------------------------------------- ACTION: ALL to add/maintain entries on http://apache.roe.ac.uk/twiki/bin/view/WFAU/ForthcomingDeadlines Continues. ACTION: MAR to investigate and quantify the issue with unmatched ATLAS sources. And if the issue warrants, RSC to look into the curation code with a view to pairing up the unmatched sources. Continues: MAR had investigated. He had reported his findings to Nigel Metcalfe and Tom Shanks and is waiting for a response from them before proceeding further. ACTION: ACD, RPB to formulate a policy for backups and for checking that data could be recovered from the backups. Continues ACTION: ACD to discuss with RGM the available budget for hardware purchases. Continues. Actions carried forward from 18/10/2013 meeting: ---------------------------------------------- ACTION NJC: RPB to check the logs, identify the error and forward to NJC, who would investigate and fix it. Continues: This relates to another VISTA non-survey release. RPB had found and forwarded the logged error to NJC. NJC now needed to retrieve the e-mail message and fix the problem. This did not get raised at the meeting so apologies if it's been discharged. ACTION: ETWS, with RPB, to develop an automated helper script to keep release DB files links updated. Continues: this one still has low priority. ACTION: NJC sort of the magnitude limits for V1.3 data. Continues ACTION: RGM will set up a doodle page to arrange the dates for remaining PDRs. Continues Specific points and new actions: ------------------------------- Project management: Nothing for the minutes. WFCAM, VISTA and VST updates: Nothing new to report. Comments and issues arising from CASU minutes: No new CASU minutes. Networking: The network had survived the UKLight upgrades and is working well. WSA/VSA/OSA/GES Operations: ETWS reported: -- Ingest of VISTA data is done for metadata ingests up to 07/13. -- Ingest of WFCAM data is done for catalogue ingests up to 07/13. -- VVV ingests are nearly done for P88. -- All the broken jpgs are reprocessed for the OSA and the VSA tiles. The deep products will finish processing over the weekend. RPB reported that he'd been looking into speeding up the VVV DR2 release. Tests had shown that the SQL Server merge command did speed up detection table outgests by a factor of 2-3. However the speed up was only achieved outgesting to the same server rather than across servers so it was not clear if this was the way forward. More tests are needed with input from RSC and NJC as required. ACD reported that the latest batch of GES data was being ingested. RGM reported that, following last weeks visit from STV and DM, most of the issues had been addressed and the new UI to the OSA would be ready for final internal checks soon. Hardware and Systems: MSH noted that RAMSES9 (VSA load server) and 10 were about to come off maintenance and other cluster nodes would follow in mid-November. ACTION: ACD/RGM to look into maintenance renewals. Webservers seshat and horus were having their Linux upgraded during the at risk period on Monday 1st November. RSC asked whether the curation servers would be upgraded to the latest Debian. MSH agreed this would be rolled out. Survey Data Release: RGM noted that Gary Davis had enquired about the final UKIDSS release. Development work was still required for items on the toDo list for the final release. An intermediate standard release was still on the cards. RGM said he'd discuss the issue with AL. VIKING - QC had been applied to the v1.3 data up to and including P90 so a release could now be started. ATLAS - as mentioned earlier, ATLAS P90 release imminent. The agreed public releases of VIKING and VMC still need to be made. Richard McMahon needed chasing up to get approval for the VHS public release. ACTION: MAR/NJC to provide RPB details of the releases to be made public. RPB to do the actual work. Software: NJC reported work on the creation of a database for 2MPZ (2MASS Photometric Redshift catalog). Initially the database would be documented and linked on the SSA website. Non-survey Data Release: No releases to report. Astrogrid deployment, VO & Data Analysis services: The ASTROTROP conference hosted by AL/IfA went well. Miscellaneous: Nothing for the minutes. -- Scanned by iCritical.